# Amortized Safe Active Learning for Real-Time Decision-Making: Pretrained Neural Policies from Simulated Nonparametric Functions

### v1

@@ 4 Experiments
Our AAL DAD GP AL AGP AL Random
1D: Airline $N_{\text{init}}+T=1+20$ $\bm{0.41\pm 0.022}$ $0.43\pm 0.018$ $0.43\pm 0.038$ $0.48\pm 0.020$ $0.42\pm 0.023$
Figure 2: Empirical results on standard AL. RMSE on airfoil dataset as a function of the query points $T$ (left). We deploy our trained policy for various query steps, i.e. $T=2,4,...,40$. For the DAD baseline, an individual policy needs to be trained for each $T$ and we train at snapshots $T=10,20,30,40$. All methods improve performance as the number of queries $T$ increases. Across all datasets, our approach is significantly faster than GP-based methods (right), as deployment involves only a single NN forward pass and no GP modeling or acquisition optimization.
Figure 3: Empirical results on safe AL. Model performance (left): Our approach (blue) performs competitively with other safe AL methods and surpasses the Safe Random baseline across all tasks. The RMSE is evaluated on safe test data only. Safety (middle): Safety awareness is quantified as the proportion of safe queries out of the total $T$ queries. Our method reaches the required safety threshold $1-\gamma$ on all datasets. Runtime (right): Compared to its competitors, our methods refrains from GP modeling and constrained acquisition optimization during querying leading to a significant speed-up. Note that Safe Random is time-consuming as well because a GP is required to estimate the safe set.
@@ 4.1 Unconstrained AL
We first study our approach on standard unconstrained AL tasks, which can be achieved by training on the objective $\mathcal{I}$ (Eq. 5) and removing the safety samples from the generative process in Algorithm 2 (see Appendix Algorithm A.4 for a cleaner version of unconstrained training). We compare our amortized AL (AAL) method against (i) GP AL: AL with a standard GP, GP trained with Type II maximum likelihood (ii) AGP AL: AL with amortized GP (AGP) proposed by Bitzer et al. (2023), where GP hyperparameters are not trained but inferred (iii) Random: random selection criterion and (iv) DAD: amortized Bayesian experimental design proposed by Foster et al. (2021) (Section C.2). We run experiments on 2 benchmark functions and 3 real-world problems: Sin function (dimension $D=1$), Branin function ($D=2$), Airline dataset ($D=1$), Langley Glide-Back Booster dataset (LGBB, $D=2$), and Airfoil dataset ($D=5$).
We report the mean and standard errors on the test dataset in Tables 1 and 2, training loss values in Appendix G. We begin by comparing the model performance: As expected, AL methods generally outperform random sampling, with the exception of the Airline dataset. Among the AL methods, we observe that the performance of AGP sometimes and of DAD often deteriorates (DAD is particularly bad when $D\geq 2$). AGP (Bitzer et al., 2023) infers hyperparameters to approximate Type II maximum likelihood, which might not fully align with the exploration needs of AL tasks. Similarly, DAD is not designed for nonparametric modeling (see Section C.2). Our AAL method, along with the DAD and Random baselines, require GPs only for evaluation purposes. In contrast, GP AL and AGP AL methods rely on GP modeling and acquisition optimization during querying, which significantly increases their runtime compared to our approach (Figure 2).
@@ 4.2 Safe AL
In this section, we study safe AL by training on the objective $\mathcal{S}_{\mathcal{H}}$ (Eq. 7) with Algorithm 2. We compare our amortized safe AL (ASAL) against the following baselines: (i) Safe GP AL (Schreiter et al., 2015; Zimmer et al., 2018; Li et al., 2022), which leverages entropy acquisition and standard GP training for the main $f$ and safety function $q$ (Eq. 1), (ii) Safe AGP AL which replaces the standard GP training by AGPs (Bitzer et al., 2023), (iii) Safe Random, where base acquisition scores are random values and a standard GP model estimates the safety values, (iv) Safe ARandom, which replaces the safety GP training by AGP. All methods are deployed with a safety level of $\gamma=0.05$.
Our method builds on the standard safe AL criteria (Eq. 1) with a focus of real-time deployment. Such criteria are usually applied to low-dimensional problems ($D\leq 2$). For higher dimensions, the acquisition optimization problem becomes highly complex and often requires approximations (Kirschner et al., 2019; Bottero et al., 2022). We deploy safe AL on two constrained benchmark functions and two real-world problems: Simionescu function ($D=2$), Townsend function ($D=2$), LGBB dataset ($D=2$) and Engine measurements ($D=3$).
Our main results are shown in Figure 3, with further analyses, including an ablation study on the effect of $\gamma$ and our novel acquisition criterion (Eq. 6) combined with conventional safe AL, presented in Appendix G. ASAL achieves comparable model performance to Safe (A)GP AL on all datasets, except for a slight disadvantage on Townsend, and outperforms Safe (A)Random across the board. It ensures safety awareness, with the proportion of safe queries exceeding the required $1-\gamma$ level on all datasets. Importantly, ASAL is at least an order of magnitude faster than all GP-based methods, including Safe (A)Random, which leverages GPs for safe set estimation. These results highlight the practical advantages of ASAL, offering a versatile solution for safe active learning in real-time applications.
@@ 5 Conclusion
We propose an amortized safe AL approach. Our approach leverages GPs to construct a simulator for pretraining the safe AL policy. The trained policy can zero-shot generalize to novel safe AL problems, enabling deployment that (i) queries informative and safe data (ii) eliminates the need for GP modeling and constrained acquisition optimization, resulting in a significant speed-up that facilitiates real-time decision making.
Many existing safe learning methods (Sui et al., 2015; Yanan Sui et al., 2018; Zimmer et al., 2018; Lederer et al., 2021) offer theoretical safety guarantees, which our method does not provide. However, these guarantees typically rely on the assumption of well-chosen GP hyperparameters a priori - an assumption that is rarely met in real-world AL applications.
Our training objectives are based on the GP likelihood, which, in its standard implementation as used here, scales cubically with dataset size. Research into linear solvers for GP training (e.g., Wang et al., 2019) offers potential alternatives to overcome this limitation. This, along with amortizable batch selection strategies (Citovsky et al., 2021), represents an active area for future research.
Future work could also extend our approach to support flexible input dimensions, as currently, a new NN policy is required for each distinct number of input features.
@@ Appendix A Gaussian Process: Distribution and Entropy
We first write down the GP predictive distribution. Details can be seen in Rasmussen & Williams (2006). We write down a general form with a non-zero prior mean $q\sim\mathcal{GP}(\mu_{q},k_{q})$. The GP hyperparameters are omitted for brevity. One can remove $\mu_{q}$ to get a zero-mean GP distribution, e.g. for $y$.
@@
$\begin{split}\log p(Z_{\text{test}}|Z)=&-n/2\log(2\pi)-1/2\log\det(cov_{\mathcal{D}}(\bm{X}_{\text{test}}))\\ &-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1/2(Z_{\text{test}}-\mu_{D}(\bm{X}_{\text{test}}))^{T}\left[cov_{\mathcal{D}}(\bm{X}_{\text{test}})\right]^{-1}(Z_{\text{test}}-\mu_{D}(\bm{X}_{\text{test}}))},\end{split}$ (A.11)
@@
$\mathbb{H}(Z_{\text{test}}|Z)=$ $\mathbb{E}_{p(Z_{\text{test}}|Z)}[-\log p(Z_{\text{test}}|Z)]=n/2\log(2\pi{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e})+1/2\log\det(cov_{\mathcal{D}}(\bm{X}_{\text{test}})).$ (A.12)
@@
where the free variables are hyperparameters of $\mu_{q},k_{q}$. The time complexity is $\mathcal{O}\left(N_{\text{observe}}^{3}\right)$ while the exact factor depends on the number of parameters, the optimizer, and the numerical stability. We keep the prior mean $\mu_{q}$ for consistent notation even though the only place when we train GPs is during deployment where we use zero mean GPs.
@@ Appendix B Policy NN Structure
Refer to caption, Refer to caption
Figure A.B.4: Our NN $\phi(T-t,\mathcal{D}_{t})=\bm{x}_{t+1}$. Each observation pair $(\bm{x}_{t},y_{t})$ (or $(\bm{x}_{t},z_{t})$) is first mapped by an MLP to an embedding $e_{t}$. Data of different $t$ use the same MLP Data Embed module, i.e. one history encoder has only one MLP Data Embed. A sequence of data embeddings is then stacked and attended by a $2$-layer transformer encoder (Vaswani et al., 2017), and then a final history embedding is obtained by summing up the attended per-data embeddings $\sum_{t}e_{t}^{a}$. The sum pool ensures an invariance of input data order. Details of the history encoder are described in Foster et al. (2021); Ivanova et al. (2021). The history encoder and safety history encoder are two modules of individual parameters. The budget variable is as well mapped by a separate MLP module to a budget embedding $EB$. The history embedding $E_{t}$, safety history embedding $ES_{t}$, and the budget embedding $EB$ are stacked, mapped by another MLP module, and then refrained by a $tahn$ function to a bounded input domain $[0,1]^{D}$.
We build our NN upon the structure described in Foster et al. (2021); Ivanova et al. (2021). The final structure is sketched in Figure A.B.4. We summarize the comparison between our network and the one developed by Ivanova et al. (2021):
@@
    the history encoder ($\{(\bm{x}_{i},y_{i})\}_{i=1}^{t}\rightarrow E_{t}$) and the decision feed forward MLP (originally $E_{t}\rightarrow\bm{x}_{t+1}$) are taken from Ivanova et al. (2021);
@@ C.1 Objectives: Unconstrained AL
0:  2.1, $T$, $N_{\text{init}}$
1:  sample a batch of $(\theta,\sigma^{2})$
2:  sample a batch of $(f,\mathcal{D}_{0})$ per Algorithm A.5
3:  for each $(f,\mathcal{D}_{0})$ do
4:     sample $T_{\text{sim}}\sim\text{Uniform}[1,T]$
5:     for $t=1,...,T_{\text{sim}}$ do
6:        $\bm{x}_{\phi,t}=\phi(T_{\text{sim}}-t+1,\mathcal{D}_{t-1})$
7:        sample $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),y_{\phi,t}=f(\bm{x}_{\phi,t})+\epsilon_{t}$
8:        $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t}\}$
9:     end for
@@
11:  compute AL loss (e.g. Eq. 5), update $\phi$
@@
0:  2.1, $N_{\text{init}}$, $\mathcal{GP}(0,k_{\theta}),\sigma^{2}$
1:  sample $f_{0}\sim\mathcal{GP}(0,k_{\theta})$
2:  $f(\cdot)=f_{0}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{0}(\bm{x})]$
3:  sample $\bm{X}_{\text{init}}\sim\text{Uniform}[\mathcal{X}],|\bm{X}_{\text{init}}|=N_{\text{init}}$
4:  sample $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I)$
5:  $Y_{\text{init}}=f(\bm{X}_{\text{init}})+\mathcal{E}_{\text{init}}$
@@ C.1.1 Objectives: Unconstrained AL - Entropy
Note that entropy $\mathbb{H}$ takes random variables, not the sampled $y_{\phi,1},...,y_{\phi,T_{\text{sim}}}$. Substituting Eqs. A.11 and A.12 into the objectives, we see that the key difference (marked blue) is indeed whether the observation values $y_{\phi,1},...,y_{\phi,T}$ are taken into account. Our main $\mathcal{H}$ aims for points $y_{\phi,1},...,y_{\phi,T}$ that are the most distinctive, while $\mathcal{H}_{\text{mean}}$ aims for $\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}$ that have the most uncertainty jointly (no actual function values). We suspect that having $y_{\phi,1},...,y_{\phi,T}$ in the loss (our main $\mathcal{H}$Eq. 4) may help the policy adapt in an AL deployment. Please see Figure A.C.5 for illustrations. One can see that $\mathcal{H}$ has the evaluated $y_{\phi,1:T_{\text{sim}}}$ which encodes actual function values and noises.
@@
Here, $y(\bm{x})$ is a random variable forecasting into the next query, while $\hat{y}(\bm{x})$ is an estimated value. In a conventional scenario where the acquisition function measures the unqueried future point, one may take the GP predictive mean for $\hat{y}(\bm{x})$ resulting in $a_{\mathcal{H}}(\bm{x}|y_{\phi,1:t},Y_{\text{init}})=-1/2\log e+\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})\propto\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})$ (Eqs. A.9, A.11 and A.12). In our simulation, $\hat{y}(\bm{x})$ would be the simulated noisy function realization (not just at $\bm{x}_{\phi,t}$ but all $\bm{x}\in\mathcal{X}$) where $\hat{y}(\bm{x}_{\phi,t+1})=y_{\phi,t+1}$.
$\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})$ is the standard predictive entropy (Seo et al., 2000; Krause & Guestrin, 2007).
We refer the readers to Krause & Guestrin (2007); Krause et al. (2008) for detailed discussions of conventional GP AL methods. In these papers, the authors discuss sequential step-wise querying and joint batch querying. Our training scheme can be considered as a special case where the policy performs sequential querying while the objectives compute joint batch querying but with true evaluations.

### v2

@@ 4 Experiments
Figure 2: Empirical results on standard AL. Left: RMSE on airfoil dataset vs number of queries $T$. Our trained policy is deployed at $T=2,4,\dots,40$. DAD trains separate policies for each $T$ (shown for $T=10,20,30,40$). All methods improve as $T$ increases. Right: total query time at $T=20$ across datasets. Our approach is significantly faster, requiring only a single NN forward pass for each data acquisition.
@@ 4.1 Unconstrained AL
Our AAL DAD GP AL AGP AL Random
1D: Airline $N_{\text{init}}+T=1+20$ $\bm{0.41\pm 0.022}$ $0.43\pm 0.018$ $0.43\pm 0.038$ $0.48\pm 0.020$ $0.41\pm 0.023$
We first study our approach on standard unconstrained AL tasks, which can be achieved by training on the objective $\mathcal{I}$ (Eq. 5) and removing the safety samples from the generative process in Algorithm 2 (see Appendix Algorithm A.4 for a cleaner version of unconstrained training). We compare our amortized AL (AAL) method against (i) GP AL: AL with a standard GP, GP trained with Type II maximum likelihood (ii) AGP AL: AL with amortized GP (AGP) proposed by Bitzer et al. [6], where GP hyperparameters are not trained but inferred (iii) Random: random selection criterion and (iv) DAD: amortized Bayesian experimental design proposed by Foster et al. [11] (Section C.2). We run experiments on 2 benchmark functions and 3 real-world problems: Sin function (dimension $D=1$), Branin function ($D=2$), Airline dataset ($D=1$), Langley Glide-Back Booster dataset (LGBB, $D=2$), and Airfoil dataset ($D=5$).
We report the mean and standard errors on the test dataset in Tables 1 and 2, training loss values in Appendix G. We begin by comparing the model performance: As expected, AL methods generally outperform random sampling, with the exception of the Airline dataset. Among the AL methods, we observe that the performance of AGP sometimes and of DAD often deteriorates (DAD is particularly bad when $D\geq 2$). AGP [6] infers hyperparameters to approximate Type II maximum likelihood, which might not fully align with the exploration needs of AL tasks. Similarly, DAD is not designed for nonparametric modeling (see Section C.2). Our AAL method, along with the DAD and Random baselines, require GPs only for evaluation purposes. In contrast, GP AL and AGP AL methods rely on GP modeling and acquisition optimization during querying, which significantly increases their runtime compared to our approach (Figure 2).
@@ 4.2 Safe AL
Figure 3: Empirical results on safe AL. Left: Our method (blue) achieves competitive RMSE and outperforms Safe Random across all tasks. The RMSE is evaluated on safe test data only. Middle: Safety awareness is quantified as the proportion of safe queries out of the total $T$ queries. Our method reaches the required threshold $1-\gamma$ on all tasks. Right: Our approach is significantly faster, as we avoid GP modeling and acquisition optimization during querying. Safe Random is also slow due to the GP-based safe set estimation.
In this section, we study safe AL by training on the objective $\mathcal{S}_{\mathcal{H}}$ (Eq. 8) with Algorithm 2. We compare our amortized safe AL (ASAL) against the following baselines: (i) Safe GP AL [40, 55, 26], which leverages entropy acquisition and standard GP training for the main $f$ and safety function $q$ (Eq. 1), (ii) Safe AGP AL which replaces the standard GP training by AGPs [6], (iii) Safe Random, where base acquisition scores are random values and a standard GP model estimates the safety values, (iv) Safe ARandom, which replaces the safety GP training by AGP. All methods are deployed with a safety level of $\gamma=0.05$.
Our method builds on the standard safe AL criteria (Eq. 1) with a focus of real-time deployment. Such criteria are usually applied to low-dimensional problems ($D\leq 3$ or $4$), as solving safe acquisitions becomes highly complex in higher dimensions where the search space expands exponentially. This often necessitates multi-stage procedures that can break the differentiability of the training objective—challenges that are beyond the scope of this work [54, 21, 7]. We deploy safe AL on two constrained benchmark functions and two real-world problems: Simionescu function ($D=2$), Townsend function ($D=2$), LGBB dataset ($D=2$) and Engine measurements ($D=3$).
Our main results are shown in Figure 3, with further analyses, including an ablation study on the effect of $\gamma$ and on our novel acquisition criterion (Eq. 7) combined with conventional safe AL, presented in Appendix G. ASAL achieves comparable model performance to Safe (A)GP AL on all datasets, except for a slight disadvantage on Townsend, and outperforms Safe (A)Random across the board. It ensures safety awareness, with the proportion of safe queries exceeding the required $1-\gamma$ level on all datasets. Importantly, ASAL is at least an order of magnitude faster than all GP-based methods, including Safe (A)Random, which leverages GPs for safe set estimation. These results highlight the practical advantages of ASAL in real-time data acquisitions.
@@ 5 Conclusion
We propose an amortized safe AL approach that leverages GPs to construct a simulator for pretraining a safe AL policy. In particular, we introduce a novel safety-aware yet unconstrained, differentiable loss and an efficient sampling scheme via GP Fourier features to simulate data acquisitions. The trained policy can zero-shot generalize to novel safe AL problems, enabling deployment that (i) queries informative and safe data, and (ii) eliminates the need for online GP modeling and constrained acquisition optimization, resulting in a significant speed-up that facilitates real-time data acquisition.
Our training relies on standard GP-based safe AL. Therefore, it inherits limitations from established GP methods and also benefits from progress in GP modeling research. A future direction here is to incorporate scalable GPs and extend to high-dimensional safe AL. Our approach is based on the idea of amortized Bayesian experimental design [11, 20, 19], where sequential acquisitions are compressed into an offline training objective. This formulation complicates the integration of multi-stage optimizations, as often required e.g. to incorporate approximate GPs (even with known GP hyperparameter samples, approximate posteriors and acquisition criteria were optimized with distinct objectives [50, 18, 54]) and higher-dimensional safe learning ($D\geq 5$[21, 7], extending beyond existing methods that typically focus on low-dimensional settings [40, 47, 4].
@@ Appendix A Gaussian process: distribution and entropy
We first write down the GP predictive distribution. Details can be seen in Rasmussen & Williams [37]. We write down a general form with a non-zero prior mean $q\sim\mathcal{GP}(\mu_{q},k_{q})$. The GP hyperparameters are omitted for brevity. One can remove $\mu_{q}$ to get a zero-mean GP distribution, e.g. for $y$.
@@
$\begin{split}\log p(Z_{\text{test}}|Z)=&-N_{\text{test}}/2\log(2\pi)-1/2\log\det(cov_{\mathcal{D}}(\bm{X}_{\text{test}}))\\ &-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1/2(Z_{\text{test}}-\mu_{D}(\bm{X}_{\text{test}}))^{T}\left[cov_{\mathcal{D}}(\bm{X}_{\text{test}})\right]^{-1}(Z_{\text{test}}-\mu_{D}(\bm{X}_{\text{test}}))},\end{split}$ (A.12)
@@
$\mathbb{H}(Z_{\text{test}}|Z)=$ $\mathbb{E}_{p(Z_{\text{test}}|Z)}[-\log p(Z_{\text{test}}|Z)]=N_{\text{test}}/2\log(2\pi{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e})+1/2\log\det(cov_{\mathcal{D}}(\bm{X}_{\text{test}})).$ (A.13)
@@
where the free variables are hyperparameters of $\mu_{q},k_{q}$ and the noise variance $\sigma^{2}$. The time complexity is $\mathcal{O}\left(N_{\text{observe}}^{3}\right)$ while the exact factor depends on the number of parameters, the optimizer, and the numerical stability. We keep the prior mean $\mu_{q}$ for consistent notation even though the only place when we train GPs is during deployment where we use zero mean GPs.
@@ Appendix B Policy NN structure
Figure A.B.4: Our NN $\phi(T-t,\mathcal{D}_{t})=\bm{x}_{t+1}$. Each observation pair $(\bm{x}_{t},y_{t})$ (or $(\bm{x}_{t},z_{t})$) is first mapped by an MLP to an embedding $e_{t}$. Data of different $t$ use the same MLP Data Embed module, i.e. one history encoder has only one MLP Data Embed. A sequence of data embeddings is then stacked and attended by a $2$-layer transformer encoder [52], and then a final history embedding is obtained by summing up the attended per-data embeddings $\sum_{t}e_{t}^{a}$. The sum pool ensures an invariance of input data order. Details of the history encoder are described in Foster et al. [11], Ivanova et al. [20]. The history encoder and safety history encoder are two modules of individual parameters. The budget variable is as well mapped by a separate MLP module to a budget embedding $EB$. The history embedding $E_{t}$, safety history embedding $ES_{t}$, and the budget embedding $EB$ are stacked, mapped by another MLP module, and then refrained by a $tahn$ function to a bounded input domain $[0,1]^{D}$.
We build our NN upon the structure described in Foster et al. [11], Ivanova et al. [20]. The final structure is sketched in Figure A.B.4. We summarize the comparison between our network and the one developed by Ivanova et al. [20]:
@@
    the history encoder ($\{(\bm{x}_{i},y_{i})\}_{i=1}^{t}\rightarrow E_{t}$) and the decision feed forward MLP (originally $E_{t}\rightarrow\bm{x}_{t+1}$) are taken from Ivanova et al. [20];
@@
Note that the history encoder incorporates the inductive bias that observed data are order-invariant (see [11, 20] for details). This can be seen by noticing that a conventional AL computes the acquisition score conditioned on the past observations and the order of the past data does not matter. We leverage the same architecture to encode the safety embedding because our safety likelihoods follows the same invariance (future likelihood conditioned on the past, see Eq. 7). We stack the history embedding, the safety history embedding, and the budget to reflect that the querying decision is made by balancing the three of them
@@ C.1 Objectives: unconstrained AL
1:2.1, $T$, $N_{\text{init}}$
2:draw a batch of $(\theta,\sigma^{2})$
3:draw a batch of $(f,\mathcal{D}_{0})$ per Algorithm A.5
4:for each $(f,\mathcal{D}_{0})$ do
5:     draw $T_{\text{sim}}\sim\text{Uniform}[1,T]$
6:     for $t=1,...,T_{\text{sim}}$ do
7:         $\bm{x}_{\phi,t}=\phi(T_{\text{sim}}-t+1,\mathcal{D}_{t-1})$
8:         draw $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),y_{\phi,t}=f(\bm{x}_{\phi,t})+\epsilon_{t}$
9:         $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t}\}$
@@
11:end for
12:compute AL loss (e.g. Eq. 5), update $\phi$
@@
1:2.1, $N_{\text{init}}$, $\mathcal{GP}(0,k_{\theta}),\sigma^{2}$
2:draw $f_{\text{raw}}\sim\mathcal{GP}(0,k_{\theta})$
3:$f(\cdot)=f_{\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{\text{raw}}(\bm{x})]$
4:draw $\bm{X}_{\text{init}}\sim\text{Uniform}[\mathcal{X}],|\bm{X}_{\text{init}}|=N_{\text{init}}$
5:draw $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I)$, $Y_{\text{init}}=f(\bm{X}_{\text{init}})+\mathcal{E}_{\text{init}}$
@@ C.1.1 Objectives: unconstrained AL - Entropy
Note that entropy $\mathbb{H}$ takes random variables, not the sampled $y_{\phi,1},...,y_{\phi,T_{\text{sim}}}$. Substituting Eqs. A.12 and A.13 into the objectives, we see that the key difference (marked blue) is indeed whether the observation values $y_{\phi,1},...,y_{\phi,T}$ are taken into account. Our main $\mathcal{H}$ aims for points $y_{\phi,1},...,y_{\phi,T}$ that are the most distinctive, while $\mathcal{H}_{\text{mean}}$ aims for $\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}$ that have the most uncertainty jointly (no actual function values). We suspect that having $y_{\phi,1},...,y_{\phi,T}$ in the loss (our main $\mathcal{H}$Eq. 4) may help the policy adapt in an AL deployment. Please see Figure A.C.5 for illustrations. One can see that $\mathcal{H}$ has the evaluated $y_{\phi,1:T_{\text{sim}}}$ which encodes actual function values and noises.
@@
Here, $y(\bm{x})$ is a random variable forecasting into the next query, while $\hat{y}(\bm{x})$ is an estimated value. In a conventional scenario where the acquisition function measures the unqueried future point, one may take the GP predictive mean for $\hat{y}(\bm{x})$ resulting in $a_{\mathcal{H}}(\bm{x}|y_{\phi,1:t},Y_{\text{init}})=-1/2\log e+\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})\propto\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})$ (Eqs. A.10, A.12 and A.13). In our simulation, $\hat{y}(\bm{x})$ would be the simulated noisy function realization (not just at $\bm{x}_{\phi,t}$ but all $\bm{x}\in\mathcal{X}$) where $\hat{y}(\bm{x}_{\phi,t+1})=y_{\phi,t+1}$.
$\mathbb{H}(y(\bm{x})|y_{\phi,1:t},Y_{\text{init}})$ is the standard predictive entropy [42, 22].
We refer the readers to Krause & Guestrin [22], Krause et al. [23] for detailed discussions of conventional GP AL methods. In these papers, the authors discuss sequential step-wise querying and joint batch querying. Our training scheme can be considered as a special case where the policy performs sequential querying while the objectives compute joint batch querying but with true evaluations.